To find the hypotenuse when side a=7, b=3, c= ? . So, a=49, b=9, c=the square root of 58 which is 7.6 according to a calculator, my question is how is the .6 derived solving this by longhand???
Answer by stanbon(75887)

To find the hypotenuse when side a=7, b=3, c= ? . So, a=49, b=9, c=the square root of 58 which is 7.6 according to a calculator, my question is how is the .6 derived solving this by longhand???
----------------
c = sqrt(49+9)
c = sqrt(58)
c = sqrt(2*29)
c = sqrt(2)*sqrt(29)
Both of these sqrt's are irrational (endless non-repeating decimals)
You can approximate the answers in decimal form but you cannot ever
completely evaluate them in decimal form.
=============================================
sqrt(58) = 7.6158.....endlessly
Cheers,
Stan H.